# Projecting Assumptions: The Duality Between Sparse Autoencoders and Concept Geometry

### v1

@@ C.2 Heterogeneity experiment
We construct a synthetic dataset consisting of five isotropic Gaussian clusters in a 128-dimensional space. The intrinsic dimensionality of each cluster follows the sequence $22^{q}-$ 2 for different values of $q\in\{3,4,5,6,7\}$, resulting in clusters with intrinsic dimensions of $6,14,30,62,126$, respectively. The lower-dimensional clusters belong to subspaces that form strict subsets of the subspaces of higher-dimensional ones, meaning that the first six dimensions are fully contained in the next 14, which are further contained in the next 30, and so on up to 126 dimensions. Cluster centers are sampled uniformly at random from the range $[0,\frac{1}{21}]$ along each dimension. The variance of each concept is chosen to be inversely proportional to its intrinsic dimension to ensure that the total variance per concept remains constant across all concepts. The dataset contains 6.4 million data points per concept, yielding a total of 32 million samples, of which 70% (approximately 22 million points) are used for training.
Our models follow four different sparse autoencoder (SAE) architectures: ReLU SAE, JumpReLU SAE, TopK SAE, and SpaDE. The first three are implemented according to their original formulations in Bricken et al. (2023), Rajamanoharan et al. (2024b), and Gao et al. (2024), with the decoder activations normalized in the forward pass. The SpaDE model follows the same single hidden-layer autoencoder structure but differs in that it utilizes Euclidean distance computations and a SparseMax activation function for the encoder. Across all models, the SAE hidden-layer width is set to 512. A pre-encoder bias is applied in all cases except for SpaDE. Additionally, for the TopK SAE, a ReLU activation is applied before selecting the top $k$ latent dimensions.
[…]
Regularization parameters are selected such that sparsity levels remain comparable across models. Specifically, the regularization coefficient $\gamma$ is chosen in the range $10^{-3}$ to $5.0$ for ReLU SAE, $10^{-3}$ to $1$ for JumpReLU SAE, from 4 to 256 (powers of 2) for TopK SAE, and from $10^{-3}$ to $10$ for SpaDE. Each model applies a different regularization strategy: ReLU SAE uses $L_{1}$ regularization, JumpReLU SAE applies $L_{0}$ regularization with a straight-through estimator (STE) following from Rajamanoharan et al. (2024b), TopK SAE does not use explicit regularization but incorporates an auxiliary loss term with $\gamma_{\text{aux}}=1$ (scaling for the auxillary term in the loss) and $K_{aux}=k$ (same as sparsity level), and SpaDE employs a distance-weighted $L_{1}$ regularization.
@@
Normalized MSE=MSEVariance of $\text{Normalized MSE}=\frac{\text{MSE}}{\text{Variance of Concept}}.$ (11)
@@ C.3 Formal Languages experiment
Data. The formal language setup analyzed in the main paper (Sec. 5.3) involves training a 2-layer nanoGPT model on strings from an English-like PCFG. Broadly, a PCFG is defined via a 5-tuple $G=(\mathtt{NT},\mathtt{T},\mathtt{R},\mathtt{S},\mathtt{P})$, where $\mathtt{NT}$ is a finite set of non-terminal symbols; $\mathtt{T}$ is a finite set of terminal symbols, disjoint from $\mathtt{NT}$; $\mathtt{R}$ is a finite set of production rules, each of the form $A\rightarrow\alpha\beta$, where $A\in\mathtt{NT}$ and $\alpha,\beta\in(\mathtt{NT}\cup\mathtt{T})$; $\mathtt{S}\in\mathtt{NT}$ is the start symbol; and $\mathtt{P}$ is a function $\mathtt{P}:\mathtt{R}\rightarrow[0,1]$, such that for each $A\in\mathtt{NT}$, $\sum_{\alpha:A\rightarrow\alpha\in\mathtt{R}}\mathtt{P}(A\rightarrow\alpha\beta)=1$. To generate a sentence from the grammar, the following process is used.
@@
We follow the same rules of the grammar considered in Menon et al. (2024). The strings are tokenized via one-hot encoding via a manually defined tokenizer.
@@
SAE training. All SAEs trained in the formal language setup involve an expansion factor of 2$\times$, i.e., 256 latents for a residual stream of 128 dimensions. Training involves a constant learning rate of $10^{-3}$ and lasts for 10K iterations ($\sim$1M tokens). We sweep regularization strength for SAEs’ training, yielding SAEs with different sparsity levels. While we fix the regularization strength for SpaDE based on best values identified from the synthetic, Gaussian cluster datasets, for other SAEs (ReLU, JumpReLU, and TopK) we report the best possible results from our sweep by looking at the top-10 per-concept F1 scores; i.e., reported results are a best-case estimate of results achievable by training of these SAEs, and in practice performance can be expected to be poorer than what we analyze. Cross-task transfer for SpaDE’s hyperparameters is intriguing in this regard, since we found other SAEs’ hyperparameters to not transfer.
@@ D.1 Projections and Nonlinearities
$z_{i}$ =$=\begin{cases}x_{i}\;\text{ if }\;x_{i}\geq 0\\ 0\;\text{else}\end{cases}$ (17)
@@
This is equivalent to ReLU, since for all non-negative values, it equals the input, while it is 00 (nearest non-negative point) for all negative inputs. Using Lemma D.1, ReLU is a projection nonlinearity with projection set $\mathcal{S}=\{\bm{x}\in\mathbb{R}^{s}:x_{i}\geq 0\forall i\in[s]\}$.
@@
$\mathbb{H}(\bm{x})_{i}=\mathbb{H}(x_{i})$ =$=\begin{cases}1\;\text{ if }x_{i}\geq 0\\ 0\;\text{ else }\end{cases}$ (23)
@@
where $s$ is the dimension of the space. Note that topK typically includes a ReLU applied first (Gao et al. (2024)), making all entries of the vector non-negative followed by choosing the $k$-largest entries of $ReLU(\bm{x})$. Consider a projection onto the union of all $k$-dimensional axis-aligned subspaces. With non-negative entries (due to ReLU), this would lead to choosing the k largest entries of $x$:
@@
[…]
@@ D.2 Receptive fields of various SAEs
[…]
@@
Given the indices =$\mathcal{M}=\{i_{1},i_{2},...,i_{|\mathcal{M}|}\}$ of active neurons (with nonzero outputs), ReLU, JumpReLU, TopK and Sparsemax are all affine functions of their inputs.
@@
$\mathcal{R}_{k}$ $=\cup_{\mathcal{M}:k\in\mathcal{M}}\mathcal{L}_{\mathcal{M}}$
@@ D.2.1 ReLU, JumpReLU SAE
$\mathcal{L}_{\mathcal{M}}$ =$=\{\bm{x}\in\mathbb{R}^{n}:\bm{w}_{m}^{T}x+b_{m}\geq 0\forall m\in\mathcal{M},\bm{w}_{q}^{T}x+b_{q}<0\forall q\notin\mathcal{M}\}$ (34)
@@
This is a consequence of the observation in Lemma D.4 and the definition of the relu model 29. ∎
@@
In Model 1 (29), for a given neuron $k\in[n]$, the receptive field $\mathcal{R}_{k}$ is given as:
$\mathcal{R}_{k}=\{\bm{x}\in\mathbb{R}^{n}:\bm{w}_{k}^{T}\bm{x}+b_{k}\geq 0\}$ (35)
@@ D.2.2 TopK SAE
$\mathcal{L}_{\mathcal{M}}$ =$=\{\bm{x}\in\mathbb{R}^{n}:\bm{w}_{m}^{T}\bm{x}\geq\bm{w}_{q}^{T}\bm{x}\forall m\in\mathcal{M},q\notin\mathcal{M}\}$ (36)
Thus, $\mathcal{M}$ is an intersection of $K(N-K)$ half-spaces all passing through the origin, and thus is a convex polytope which may be bounded or unbounded. In fact, it is an unbounded hyperpyramid, with a corner at the origin and flat faces. The normals to these half-spaces are pairwise differences between active and inactive weight vectors.
@@
In Model 2 (31), for a given neuron $k\in[n]$, the receptive field $\mathcal{R}_{k}$ is given as:
$\mathcal{R}_{k}=\cup_{\mathcal{M}:k\in\mathcal{M}}\mathcal{L}_{\mathcal{M}}$ (37)
@@ D.2.3 SpaDE
$\mathcal{L}_{\mathcal{M}}$ =$=\Bigg{\{}\bm{x}\in\mathbb{R}^{n}:\|\bm{x}-\bm{w}_{m}\|_{2}^{2}-\frac{1}{|\mathcal{M}|}\sum_{j\in\mathcal{M}}\|\bm{x}-\bm{w}_{j}\|_{2}^{2}-\frac{1}{\lambda|\mathcal{M}|}\;\begin{cases}\leq 0,\;\;\text{if }m\in\mathcal{M}\\ >0,\;\;m\notin\mathcal{M}\end{cases}\Bigg{\}}$ (38)
=$:\displaystyle=\Bigg{\{}\bm{x}\in\mathbb{R}^{n}$ : (39)
@@
This is again a consequence of the definition of sparsemax Martins & Astudillo (2016). ∎
@@
In SpaDE (32), for a given neuron $k\in[n]$, the receptive field $\mathcal{R}_{k}$ is given as:
$\mathcal{R}_{k}=\cup_{\mathcal{M}:k\in\mathcal{M}}\mathcal{L}_{\mathcal{M}}$ (41)
@@ D.3 KDS and Sparse Coding
K-Deep Simplex (KDS) Tasissa et al. (2023) is the sparse coding framework which forms the outer optimization in the SpaDE. While this is a different framework, in this section we show that it is general enough to capture the standard sparse coding, i.e., for data generated using standard sparse coding, there exists a corresponding KDS framework that could have generated the same data. Note that we may have to increase the latent dimension (number of dictionary atoms) by one to obtain the corresponding KDS framework. This is stated and proved (with a constructive proof) in the following theorem.
@@
Given data $\mathcal{D}=\{\bm{x}^{(1)},...,\bm{x}^{(P)}\}$ generated from a standard sparse coding generative model, i.e., $\bm{x}=\bm{D}\bm{z}+\eta$, where dictionary atoms (columns of $\bm{D}$) have unit norm and $\bm{z}$ is unconstrained, there exists a scaling of the data such that it can be represented using the K-Deep Simplex Tasissa et al. (2023) framework, i.e., $\tilde{\bm{x}}=\kappa\bm{x}=\tilde{\bm{D}}\tilde{\bm{z}}+\tilde{\eta}$, where $\tilde{\bm{z}}\in\Delta^{s}$.
@@ D.4 SpaDE
Sparsemax is a projection onto the probability simplex, which can be written as (see Proposition 1 in Martins & Astudillo (2016))
Let $\text{Let }\bm{z}$ $=\text{Sparsemax}(\bm{y})$ (42)
[…]
@@
$\implies\text{Sparsemax}(\bm{y})_{i}$ $=\text{ReLU}\left(2\lambda(\bm{w}_{i}-\frac{1}{|\mathcal{M}|}\sum_{j}\bm{w}_{j})^{T}\bm{x}-\lambda(|\bm{w}_{i}|^{2}-\frac{1}{|\mathcal{M}|}|\bm{w}_{j}|^{2})+\frac{1}{|\mathcal{M}|}\right)$ (45)
@@
where $\tilde{\bm{W}}(\bm{x})=2\lambda(\bm{w}_{i}-\frac{1}{|\mathcal{M}|}\sum_{j}\bm{w}_{j})$, $\tilde{\bm{b}}_{e}=-\lambda(|\bm{w}_{i}|^{2}-\frac{1}{|\mathcal{M}|}|\bm{w}_{j}|^{2})+\frac{1}{|\mathcal{M}|}$ and $\mathcal{M}$ is the set of active indices, which is uniquely determined by the constraint $\sum_{i}\text{Sparsemax}(\bm{y})_{i}=1$ (see Proposition 1 in Martins & Astudillo (2016) for uniqueness). Note that $\tilde{\bm{W}}(\bm{x}),\tilde{\bm{b}}_{e}(\bm{x})$ are both piecewise constant on regions of input space marked by the same choice of active indices.
@@
This resembles one-sided optimal transport with a squared 2-norm regularizer. This problem is one-sided because there is no constraint on how much weight sits on each prototype across different inputs (optimization is performed independently for each input). The squared $2-$norm regularizer is known to lead to sparse transport plans in the optimal transport literature (see Liu et al. (2022)).
The outer optimization for SpaDE (Eq. 4) is a locality-enforced version of dictionary learning called K-Deep Simplex (KDS) Tasissa et al. (2023). In this framework, the sparse code is constrained to belong to the probability simplex, i.e., $\bm{z}\in\Delta^{s}=\{\bm{y}\in\mathbb{R}^{s}:\sum_{i}y_{i}=1,y_{i}\geq 0\forall i\}$, while the dictionary atoms $\bm{D}$ are unconstrained. The distance-weighted L1 regularizer encourages each datapoint to use those dictionary atoms which are close to itself in euclidean distance, inducing a soft clustering bias. Even though this is a different dictionary learning framework than standard sparse coding, it is expressive enough to capture the standard sparse coding setup, i.e., for any standard sparse coding problem, there exists an equivalent KDS problem (see Theorem D.8 in Appendix).
@@
    Constraining the sparse code to the probability simplex finds support in an oft-cited paper demonstrating the linear representation hypothesis in word embeddings under a random-walk based generative model of language Arora et al. (2018). Their main result (Theorem 2) shows that representations are convex combinations of concepts, as opposed to unconstrained linear combinations, which is better interpreted as assigning vectors (with magnitude and direction; alternatively, locations) to concepts rather than directions (without magnitude). This idea of concepts as vectors has also been demonstrated both theoretically and empirically in the final layer representations of language models Park et al. (2024).
@@
    The projection set $\mathcal{S}$ in SpaDE is the probability simplex, which admits edges/corners with varying levels of sparsity, thereby allowing the representation of heterogeneous concepts.

### v2

@@ C.2 Heterogeneity experiment
We construct a synthetic dataset consisting of five isotropic Gaussian clusters in a 128-dimensional space. The intrinsic dimensionality of each cluster follows the sequence $2^{q}-2$ for different values of $q\in\{3,4,5,6,7\}$, resulting in clusters with intrinsic dimensions of $6,14,30,62,126$, respectively. The lower-dimensional clusters belong to subspaces that form strict subsets of the subspaces of higher-dimensional ones, meaning that the first six dimensions are fully contained in the next 14, which are further contained in the next 30, and so on up to 126 dimensions. Cluster centers are sampled uniformly at random from the range $0,\frac{1}{21}]$ along each dimension. The variance of each concept is chosen to be inversely proportional to its intrinsic dimension to ensure that the total variance per concept remains constant across all concepts. The dataset contains 6.4 million data points per concept, yielding a total of 32 million samples, of which 70% (approximately 22 million points) are used for training.
Our models follow four different sparse autoencoder (SAE) architectures: ReLU SAE, JumpReLU SAE, TopK SAE, and SpaDE. The first three are implemented according to their original formulations in [12], [14], and [13], with the decoder activations normalized in the forward pass. The SpaDE model follows the same single hidden-layer autoencoder structure but differs in that it utilizes Euclidean distance computations and a SparseMax activation function for the encoder. Across all models, the SAE hidden-layer width is set to 512. A pre-encoder bias is applied in all cases except for SpaDE. Additionally, for the TopK SAE, a ReLU activation is applied before selecting the top $k$ latent dimensions.
For training, the temperature parameter $\lambda$ in SpaDE is initialized at $1/(2\times\text{input dimension})$ and parameterized using the Softplus function to ensure non-negativity. This parameter trained along with the encoder and decoder weights, to allow the model to learn its desired sparsity level. In JumpReLU, the threshold is initialized at $10^{-3}$ across all latent dimensions, with a bandwidth of $10^{-3}$ for the straight-through estimator (STE). All models are trained using the Adam optimizer with a learning rate of $10^{-2}$, which follows a cosine decay schedule from $10^{-2}$ to $10^{-4}$. The momentum parameter is set to 0.9, and we use a batch size of 2048. Training runs for approximately 10,000 iterations, and gradient clipping (restricting gradient norms to be less than 1) is applied to stabilize optimization.
Regularization parameters are selected such that sparsity levels remain comparable across models. Specifically, the regularization coefficient $\gamma$ is chosen in the range $10^{-3}$ to $5.0$ for ReLU SAE, $10^{-3}$ to $1$ for JumpReLU SAE, from 4 to 256 (powers of 2) for TopK SAE, and from $10^{-3}$ to $10$ for SpaDE. Each model applies a different regularization strategy: ReLU SAE uses $L_{1}$ regularization, JumpReLU SAE applies $L_{0}$ regularization with a straight-through estimator (STE) following from [14], TopK SAE does not use explicit regularization but incorporates an auxiliary loss term with $\gamma_{\text{aux}}=1$ (scaling for the auxillary term in the loss) and $K_{aux}=k$ (same as sparsity level), and SpaDE employs a distance-weighted $L_{1}$ regularization.
@@
$\text{Normalized MSE}=\frac{\text{MSE}}{\text{Variance of Concept}}.$ (11)
@@ C.3 Formal Languages experiment
Data. The formal language setup analyzed in the main paper (Sec. 5.3) involves training a 2-layer nanoGPT model on strings from an English-like PCFG (Probabilistic Context-Free Grammars). Broadly, a PCFG is defined via a 5-tuple $G=(\mathtt{NT},\mathtt{T},\mathtt{R},\mathtt{S},\mathtt{P})$, where $\mathtt{NT}$ is a finite set of non-terminal symbols; $\mathtt{T}$ is a finite set of terminal symbols, disjoint from $\mathtt{NT}$; $\mathtt{R}$ is a finite set of production rules, each of the form $A\rightarrow\alpha\beta$, where $A\in\mathtt{NT}$ and $\alpha,\beta\in(\mathtt{NT}\cup\mathtt{T})$; $\mathtt{S}\in\mathtt{NT}$ is the start symbol; and $\mathtt{P}$ is a function $\mathtt{P}:\mathtt{R}\rightarrow[0,1]$, such that for each $A\in\mathtt{NT}$, $\sum_{\alpha:A\rightarrow\alpha\in\mathtt{R}}\mathtt{P}(A\rightarrow\alpha\beta)=1$. To generate a sentence from the grammar, the following process is used.
@@
We follow the same rules of the grammar considered in [34]. The strings are tokenized via one-hot encoding via a manually defined tokenizer.
@@
[…]
@@ D.1 Projections and Nonlinearities
The nonlinearity of popular SAEs is commonly an orthogonal projection onto some set, where the choice of projection set differentiates SAEs (see Fig. 2). We formalize such nonlinearities as projection nonlinearities, as (re)defined below.
Definition D.1 (Projection Nonlinearity).
Model $\bm{g}(\bm{v})$
[…]
We will say a function $\bm{f}(\cdot)$ is a Projection Encoder if it uses a projection nonlinearity $\bm{g}(\cdot)$ applied to a linear transformation of the input. This is equivalent to using $\bm{v}=\bm{W}^{\mathsf{T}}\bm{x}+\bm{b}_{e}$, and $\bm{f}=\bm{g}(\bm{v})$ (see Eq. 2), where $\bm{g}$ is a projection nonlinearity. Popular SAEs can be understood as a similar Projection Encoder with different projection nonlinearities, as shown in Tab. 4 (see Theorem D.3 for a derivation).
@@
$z_{i}$ $=\begin{cases}x_{i}\;\text{ if }\;x_{i}\geq 0\\ 0\;\text{else}\end{cases}$ (18)
@@
This is equivalent to ReLU, since for all non-negative values, it equals the input, while it is 0 (nearest non-negative point) for all negative inputs. Using Lemma D.2, ReLU is a projection nonlinearity with projection set $\mathcal{S}=\{\bm{x}\in\mathbb{R}^{s}:x_{i}\geq 0\forall i\in[s]\}$.
@@
[…]
@@
where $s$ is the dimension of the space. Note that topK typically includes a ReLU applied first ([13]), making all entries of the vector non-negative followed by choosing the $k$-largest entries of $ReLU(\bm{x})$. Consider a projection onto the union of all $k$-dimensional axis-aligned subspaces. With non-negative entries (due to ReLU), this would lead to choosing the k largest entries of $x$:
@@
    $\bm{\Pi}_{\mathcal{S}}\left\{\bm{x}\right\}$ $=\bm{c},\;\bm{x}\in\bm{D}\text{ (a linear piece)}$
@@ D.2 Receptive fields of various SAEs
$\text{ReLU SAE: }\;\bm{z}$ $=\text{ReLU}(\bm{W}^{T}\bm{x}+\bm{b})$ (30)
$\text{JumpReLU SAE: }\;\bm{z}$ $=\text{JumpReLU}(\bm{W}^{T}\bm{x}+\bm{b})$ (31)
$\text{TopK SAE: }\;\bm{z}$ $=\text{TopK}(\bm{W}^{T}\bm{x})$ (32)
$\text{SpaDE: }\;\bm{z}$ $=\text{Sparsemax}(-\lambda d(\bm{x},\bm{W}))$ (33)
@@
Intuition behind Theorem 4.1
Receptive fields for ReLU and JumpReLU SAEs are half-spaces because of the linear transform in the encoder, which needs to be positive for the SAE latent to be active (due to the ReLU/ JumpReLU nonlinearity). In TopK SAE, for a given latent to be active, its linear transform in the encoder must be non-negative (since a ReLU is used in TopK SAE implementation [13]), and must exceed at least sKs-K other linear transforms (where $s$ is the SAE width). This gives us an intersection of multiple half-spaces through the origin, leading to hyperpyramids and thus angular separation.
For projection-based encoders, the receptive field can be rewritten as
$\mathcal{F}_{k}=\bm{f}^{-1}\big(\mathcal{S}\cap\{z_{k}>0\}\big),$
where $\mathcal{S}$ is the projection set of the encoder.
That is, $\mathcal{F}_{k}$ is the pre-image of the intersection of the projection set with the half-space $\{z_{k}>0\}$. Alternatively, it can be viewed as the complement of the pre-image of the set $\mathcal{S}\cap\{z_{k}=0\}$, where the hyperplane $z_{k}=0$ indicates latent $k$ is “dead”. This expression shows the explicit relation between the projection set and the receptive field properties of the SAE.
@@
Given the indices $\mathcal{M}=\{i_{1},i_{2},...,i_{|\mathcal{M}|}\}$ of active neurons (with nonzero outputs), ReLU, JumpReLU, TopK and Sparsemax are all affine functions of their inputs.
@@
$\mathcal{F}_{k}$ $=\cup_{\mathcal{M}:k\in\mathcal{M}}\mathcal{L}_{\mathcal{M}}$
@@ D.2.1 ReLU, JumpReLU SAE
$\mathcal{L}_{\mathcal{M}}$ $=\{\bm{x}\in\mathbb{R}^{n}:\bm{w}_{m}^{T}x+b_{m}\geq 0\forall m\in\mathcal{M},\bm{w}_{q}^{T}x+b_{q}<0\forall q\notin\mathcal{M}\}$ (35)
@@
This is a consequence of the observation in Lemma D.5 and the definition of the relu model 30. ∎
@@
In Model 1 (30), for a given neuron $k\in[n]$, the receptive field $\mathcal{F}_{k}$ is given as:
$\mathcal{F}_{k}=\{\bm{x}\in\mathbb{R}^{n}:\bm{w}_{k}^{T}\bm{x}+b_{k}\geq 0\}$ (36)
@@ D.2.2 TopK SAE
$\mathcal{L}_{\mathcal{M}}$ $=\{\bm{x}\in\mathbb{R}^{n}:\bm{w}_{m}^{T}\bm{x}\geq\bm{w}_{q}^{T}\bm{x}\forall m\in\mathcal{M},q\notin\mathcal{M}\}$ (37)
Thus, $\mathcal{M}$ is an intersection of K(NK)K(N-K) half-spaces all passing through the origin, and thus is a convex polytope which may be bounded or unbounded. In fact, it is an unbounded hyperpyramid, with a corner at the origin and flat faces. The normals to these half-spaces are pairwise differences between active and inactive weight vectors.
@@
In Model 2 (32), for a given neuron $k\in[n]$, the receptive field $\mathcal{F}_{k}$ is given as:
$\mathcal{F}_{k}=\cup_{\mathcal{M}:k\in\mathcal{M}}\mathcal{L}_{\mathcal{M}}$ (38)
@@ D.2.3 SpaDE
$\mathcal{L}_{\mathcal{M}}$ $=\Bigg\{\bm{x}\in\mathbb{R}^{n}:\|\bm{x}-\bm{w}_{m}\|_{2}^{2}-\frac{1}{|\mathcal{M}|}\sum_{j\in\mathcal{M}}\|\bm{x}-\bm{w}_{j}\|_{2}^{2}-\frac{1}{\lambda|\mathcal{M}|}\;\begin{cases}\leq 0,\;\;\text{if }m\in\mathcal{M}\\ >0,\;\;m\notin\mathcal{M}\end{cases}\Bigg\}$ (39)
$=\Bigg\{\bm{x}\in\mathbb{R}^{n}:$ (40)
@@
This is again a consequence of the definition of sparsemax [44]. ∎
@@
In SpaDE (33), for a given neuron $k\in[n]$, the receptive field $\mathcal{F}_{k}$ is given as:
$\mathcal{F}_{k}=\cup_{\mathcal{M}:k\in\mathcal{M}}\mathcal{L}_{\mathcal{M}}$ (42)
@@ D.3 KDS and Sparse Coding
K-Deep Simplex (KDS) [45] is the sparse coding framework which forms the outer optimization in the SpaDE. While this is a different framework, in this section we show that it is general enough to capture the standard sparse coding, i.e., for data generated using standard sparse coding, there exists a corresponding KDS framework that could have generated the same data. Note that we may have to increase the latent dimension (number of dictionary atoms) by one to obtain the corresponding KDS framework. This is stated and proved (with a constructive proof) in the following theorem.
@@
Given data $\mathcal{D}=\{\bm{x}^{(1)},...,\bm{x}^{(P)}\}$ generated from a standard sparse coding generative model, i.e., $\bm{x}=\bm{D}\bm{z}+\eta$, where dictionary atoms (columns of $\bm{D}$) have unit norm and $\bm{z}$ is unconstrained, there exists a scaling of the data such that it can be represented using the K-Deep Simplex [45] framework, i.e., $\tilde{\bm{x}}=\kappa\bm{x}=\tilde{\bm{D}}\tilde{\bm{z}}+\tilde{\eta}$, where $\tilde{\bm{z}}\in\Delta^{s}$.
@@ D.4 SpaDE
Sparsemax is a projection onto the probability simplex, which can be written as (see Proposition 1 in [44])
$\text{Let }\bm{z}$ $=\text{Sparsemax}(\bm{y})$ (43)
$\text{Then, }z_{i}$ $=\text{ReLU}(y_{i}-\frac{1}{|\mathcal{M}|}\sum_{j\in\mathcal{M}}y_{j}+\frac{1}{|\mathcal{M}|})$ (44)
@@
$\implies\text{Sparsemax}(\bm{y})_{i}$ $=\text{ReLU}\left(2\lambda(\bm{w}_{i}-\frac{1}{|\mathcal{M}|}\sum_{j}\bm{w}_{j})^{T}\bm{x}-\lambda(|\bm{w}_{i}|^{2}-\frac{1}{|\mathcal{M}|}\sum_{j}|\bm{w}_{j}|^{2})+\frac{1}{|\mathcal{M}|}\right)$ (46)
@@
where $\tilde{\bm{W}}(\bm{x})=2\lambda(\bm{w}_{i}-\frac{1}{|\mathcal{M}|}\sum_{j}\bm{w}_{j})$, $\tilde{\bm{b}}_{e}=-\lambda(|\bm{w}_{i}|^{2}-\frac{1}{|\mathcal{M}|}\sum_{j}|\bm{w}_{j}|^{2})+\frac{1}{|\mathcal{M}|}$ and $\mathcal{M}$ is the set of active indices, which is uniquely determined by the constraint $\sum_{i}\text{Sparsemax}(\bm{y})_{i}=1$ (see Proposition 1 in [44] for uniqueness). Note that $\tilde{\bm{W}}(\bm{x}),\tilde{\bm{b}}_{e}(\bm{x})$ are both piecewise constant on regions of input space marked by the same choice of active indices.
@@
However, SpaDE differs from a ReLU SAE by using linear transformations defined with respect to a local origin, which is uniquely determined by the set of active SAE latents, similar to recent work on steering [80].
@@
This resembles one-sided optimal transport with a squared 2-norm regularizer. This problem is one-sided because there is no constraint on how much weight sits on each prototype across different inputs (optimization is performed independently for each input). The squared 22-norm regularizer is known to lead to sparse transport plans in the optimal transport literature (see [81]).
The outer optimization for SpaDE (Eq. 4) is a locality-enforced version of dictionary learning called K-Deep Simplex (KDS) [45]. In this framework, the sparse code is constrained to belong to the probability simplex, i.e., $\bm{z}\in\Delta^{s}=\{\bm{y}\in\mathbb{R}^{s}:\sum_{i}y_{i}=1,y_{i}\geq 0\forall i\}$, while the dictionary atoms $\bm{D}$ are unconstrained. The distance-weighted L1 regularizer encourages each datapoint to use those dictionary atoms which are close to itself in euclidean distance, inducing a soft clustering bias. Even though this is a different dictionary learning framework than standard sparse coding, it is expressive enough to capture the standard sparse coding setup, i.e., for any standard sparse coding problem, there exists an equivalent KDS problem (see Theorem D.9 in Appendix).
@@
    Constraining the sparse code to the probability simplex finds support in an oft-cited paper demonstrating the linear representation hypothesis in word embeddings under a random-walk based generative model of language [9]. Their main result (Theorem 2) shows that representations are convex combinations of concepts, as opposed to unconstrained linear combinations, which is better interpreted as assigning vectors (with magnitude and direction; alternatively, locations) to concepts rather than directions (without magnitude). This idea of concepts as vectors has also been demonstrated both theoretically and empirically in the final layer representations of language models [82].
@@
    The projection set $\mathcal{S}$ in SpaDE is the probability simplex, which admits edges/corners with varying levels of sparsity, thereby allowing the representation of heterogeneous concepts. For any choice of $k\in\{1,2,...,s\}$, there are $s\choose k$ choices of indices $\mathcal{M}_{k}$ for a $k$-sparse representation, and points $\{\bm{x}\in\mathbb{R}^{s}:x_{i}=0,i\notin\mathcal{M}_{k},\sum_{j\in\mathcal{M}_{k}}x_{j}=1,x_{j}\geq 0\}\subseteq\Delta^{s}$ which admit this level of sparsity, thereby capturing concept heterogeneity.